OpenStudy (anonymous):

If a and b are positive numbers, find the maximum value of f(x)=x^a(1-x)^b, 0<=x<=1

OpenStudy (anonymous):

\[0\le x \le1\]

OpenStudy (anonymous):

\[x^a(1-x)^b\]

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

finding maximum denotes getting the 1st derivative of the function and equate to zero...

OpenStudy (anonymous):

ok so the derivative would be ax^(a-1)(1-x)^b-b(1-x)^(b-1)(x^a)

OpenStudy (anonymous):

isn't it + since that d (uv) = u dv + v du

OpenStudy (anonymous):

yeah but with the chain rule you have to take the derivative of (1-x) which is -1 and multiply it to the exponent rule part so it becomes -b

OpenStudy (anonymous):

oh yes i overlook (1-x) (-1) dx...

OpenStudy (anonymous):

so what do i do now

OpenStudy (anonymous):

combine like terms, arrange in d[f(x)]/dx then equate to 0 to solve for x in terms of a and b...

OpenStudy (anonymous):

equating to zero (0) the 1st derivative denotes either maximum or minimum....

OpenStudy (anonymous):

i have b(1-x)^b-1(x^a)=ax^(a-1)(1-x)^b

OpenStudy (anonymous):

\[x^a(1-x)^b=>bx^a(1-x)^{b-1}+a(1-x)^b x^{a-1}=0\] solve for x=?

OpenStudy (anonymous):

it should be -

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

\[bx^a(1-x)^{b-1}=a(1-x)^b x^{a-1}\]

OpenStudy (anonymous):

oops.. we can cancel (1-x)^b on both sides...

OpenStudy (anonymous):

exactly^ :D sorry im eating so didnt type..was waiting fr sm1 to observe that!

OpenStudy (anonymous):

but its (1-x)^b-1 on the left

OpenStudy (anonymous):

\[\frac{bx ^{a}}{(1-x)}=ax ^{a-1}\]

OpenStudy (anonymous):

we can further cancel (x^a)...

OpenStudy (anonymous):

\[\LARGE bx^a(1-x)^{b-1}=a(1-x)^b x^{a-1}\] \[\frac{bx^a \cancel{(1-x)^{b}}}{(1-x)}=a \cancel{(1-x)^b} x^{a-1}\]

OpenStudy (anonymous):

\[\frac{ b }{ 1-x }=\frac{ a }{ x }\]

OpenStudy (anonymous):

solve for x=0..and solve for maxima minima

OpenStudy (anonymous):

if x=0 then the left side of @Orion1213 equation is undefined

OpenStudy (anonymous):

bx = a - ax bx + ax = a (a+b)x = a x = a/(a+b)

OpenStudy (anonymous):

did u get it @muzzammil.raza

OpenStudy (anonymous):

i got lost after orion said b/1-x=a/x and then you said x=0 which makes the equation undefined

OpenStudy (anonymous):

we have solve x here, the condition must be satisfied that a and b are both positive numbers...

OpenStudy (anonymous):

and maximum value of f(x) occurred at x=a/(a+b)...

OpenStudy (anonymous):

... regarding the condition for x.... 0<=x<=1... if a is positive, it should be greater than 0 so x is safe of not to be zero....

OpenStudy (anonymous):

but how did you get from b/1-x=a/x to bx=a-ax

OpenStudy (anonymous):

if we go back to... \[\frac{ bx ^{a} }{ 1-x }=ax ^{a-1}\]

OpenStudy (anonymous):

we got this form when we cancel (1-x)^b...

OpenStudy (anonymous):

yes i get that part

OpenStudy (anonymous):

examining further the equation, we can further cancel x^a...

OpenStudy (anonymous):

note on the right-side... \[x ^{a-1}=x ^{a}\times x ^{-1}=\frac{ x ^{a} }{ x }\]

OpenStudy (anonymous):

so we can re-write the equation as... \[\frac{ bx ^{a} }{ 1-x }=\frac{ ax ^{a} }{ x }\]

OpenStudy (anonymous):

if we multiply both sides by 1/(x^a)... x^a will cancel and leaves the equation as \[\frac{ b }{ 1-x }=\frac{ a }{ x }\]

OpenStudy (anonymous):

from this form... we can now solve for the value of x... multiply both sides by x(1-x).... \[x \cancel{(1-x)}\left[ \frac{ b }{ \cancel{(1-x)} } \right]=\left[ \frac{ a }{ \cancel{x} } \right](1-x)\cancel{x}\]

OpenStudy (anonymous):

bx = a (1-x) bx = a - ax ax + bx = a (a+b)x = a \[x=\frac{ a }{ a+b }\]

OpenStudy (anonymous):

@muzzammil.raza , is it ok now... got the idea?

OpenStudy (anonymous):

so the maximum is a/a+b

OpenStudy (anonymous):

not yet... the requirement is the maximum of f(x)... so we need to substitute x in our function... and that is the answer we are looking for....

OpenStudy (anonymous):

\[f(x) = x ^{a}(1-x)^{b}\]

OpenStudy (anonymous):

ok so the maximum is (a/(a+b))^a(1-(a/a+b))^b

OpenStudy (anonymous):

yes... that is the one we're looking for... :)

OpenStudy (anonymous):

can i simplify that

OpenStudy (anonymous):

it seems complicated to me already... :)

OpenStudy (anonymous):

is that a yes or no

OpenStudy (anonymous):

for me... no...

OpenStudy (anonymous):

ok i will leave it then thanks for all the help

OpenStudy (anonymous):

no problem... :)